OpenStudy (anonymous):

Solve 15x2 + 23x + 20 = 16 by factoring.

OpenStudy (anonymous):

15x2 + 23x+4=0 15x2+20x+3x+4=0 5x(3x+4)+1(3x+4)=0 (3x+4)(5x+1)=0 x=-4/3 x=-1/5